OpenStudy (anonymous):

Write the equation of the line that passes through the points (3, 6) and (4, 10) using function notation: A. F(x) = 4x - 6 B. F(x) = x + 4 C. Y = x + 4 D. Y = 4x - 6

OpenStudy (anonymous):

I know it's not c and d but I don't know if it's a or b

ganeshie8 (ganeshie8):

plugin the points and see which option satisifes both the points

ganeshie8 (ganeshie8):

(3, 6) plugin x = 3 in first option : F(3) = 4*3 - 6 = 12 - 6 = 6 so (3, 6) is indeed on this line similarly test the other point also

OpenStudy (anonymous):

How would I do that? @ganeshie8

OpenStudy (zzr0ck3r):

Or you can actually figure out the equation of a line containing those two points

OpenStudy (anonymous):

So is it A @ganeshie8

ganeshie8 (ganeshie8):

how did you get A ?

OpenStudy (anonymous):

So would it be b bc the question is asking for a function notation @ganeshie8

OpenStudy (usukidoll):

function notation is f(x) so the last 2 choices are knocked out already.

OpenStudy (usukidoll):

I think @ganeshie8 wants you to show your work on you got A as the answer.